in a cyclic quadrilateral angle A is 20 digree grater than C.then find angle AandC
Answer by jim_thompson5910(35256)

A = C + 20

A+C = 180

C+20+C = 180

2C+20 = 180

2C = 180 - 20

2C = 160

C = 160/2

C = 80

A = C + 20

A = 80 + 20

A = 100

Answers:

A = 100
C = 80